A town house can be rented for P10,000 for three days, but for
the fourth day onwards, an additional cost of P3,000 per day is
added. Represent the cost of renting the town house as a
piecewise function of the number of days it is rented.

Found 2 solutions by Theo, MathTherapy:
Answer by Theo(13342)

y = 10,000 + 0 * x for x = 1 to 3
y = 10000 + 3000 * (x-3) for x > 3

y represents the total cost.
x represents the number of days.

the graph would look something like this.



the total cost is 10,000 for days 1 to 3.
it is 10,000 + 3000 * (x-3) for days > 3.
in the graph, day 6 is shown.
the total cost for day 6 is 10,000 + 3 * 3000 = 19,000.







Answer by MathTherapy(10556)

A town house can be rented for P10,000 for three days, but for
the fourth day onwards, an additional cost of P3,000 per day is
added. Represent the cost of renting the town house as a
piecewise function of the number of days it is rented.
With cost being C, and number of days, d, we get: